A d.c. motor has the following spec: k×flux= 6.0v.s Vt= 10:0 V Ra= 1.5:00 Ia= 4:0 A (armature current at full load). Calc rated torque, starting torque; starting current at full voltage, starting speed; For Seperately excited and Shunt motor?
1
Expert's answer
2020-06-10T07:49:44-0400

As per the given in the question,

k×flux= 6.0v.s

Vt= 10:0 V

Ra= 1.5:00

Ia= 4:0 A

We know that rated torque "(T)=k\times flux\times I_a"

"=6\times 4 =24Vs A"

suppose k =1

Starting torque "=KI_a^2" "=4^2=16VsA"

starting current "(I)=\frac{Vt}{R}=\frac{10}{1.5}=6.66A"

starting torque for the shunt motor = KIa= 4VsA